OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -4, -16, -64, -256, ...

OpenStudy (anonymous):

an = a1 + (n – 1)d

OpenStudy (usukidoll):

I think we need to just find an equation that produces that string

OpenStudy (usukidoll):

-(4)^n and let n = 1,2,3,4...

OpenStudy (zzr0ck3r):

-(4)^1 = -4 -(4)^2 = -16 -(4)^3 = ?

OpenStudy (usukidoll):

lol ^ that's what I'm thinking too

OpenStudy (anonymous):

-64

OpenStudy (usukidoll):

think of it as this... ok there are - signs everywhere so we really need -... next I see a 4 16 64... hmmm square root of 16 is 4... 4^ 2 = 16.. cube root of 64 is 4... 4^3 = 64

OpenStudy (anonymous):

an = 4 • -4n an = 4 • -4n + 1 an = -4 • 4n an = -4 • 4n - 1

OpenStudy (usukidoll):

so your equation would be -(4)^n

OpenStudy (usukidoll):

we don't need a_n

OpenStudy (usukidoll):

-(4)^n is the answer.. what if we let n = 1 2 3 4? then we have -4,-16,-64,-256 which is what you want right?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

^^ then that's the answer -(4)^n is your equation

OpenStudy (anonymous):

= -4 • 4n?

OpenStudy (usukidoll):

no ^

OpenStudy (anonymous):

4 • -4n

OpenStudy (usukidoll):

ignore JoyLin's suggestion... that's wrong

OpenStudy (anonymous):

because -4n is not one of my answer choices

OpenStudy (anonymous):

an = 4 • -4n an = 4 • -4n + 1 an = -4 • 4n an = -4 • 4n - 1

OpenStudy (anonymous):

thats my choices

OpenStudy (usukidoll):

oh strange ... because you could just use what zzrocker and I said and that would be fine

OpenStudy (anonymous):

Thank you c: